OpenStudy (anonymous):

How to determine the magnetic field due to a straight current-carrying conductor of finite size?

OpenStudy (vincent-lyon.fr):

Integrate Biot & Savart law.

OpenStudy (anonymous):

The magnetic field due dB due to an element dL carrying a steady current I at a point P at a distance r from the current element is \[dB=(\mu0/4\pi) I *((dL*r)/r ^{3}))\]

OpenStudy (anonymous):

How about a conductor of finite size?

OpenStudy (anonymous):

Integrate this vector expresion

OpenStudy (anonymous):

There is another one...The magnitude of the mag field at a distance R from a long straight wire carrying a current I is given by \[B=\mu0I/2\Pi r\]

OpenStudy (anonymous):

But this is the result for a wire of infinite length..

OpenStudy (anonymous):

I am asking for finite length..

OpenStudy (anonymous):

 then magnetic field at point p is given by\[B = \mu o I/4 \pi r ( \sin a + \sin b )\]